A firm produces its output using only capital and labor.Labor costs $100 per worker per day and capital costs $200 per unit per day. If the marginal product of the last worker employed is 500 and the marginal product of the last unit of capital is 1,000, is the firm employing the cost-minimizing combination of inputs? Explain.
Answer #1

Cost is minimized when (Marginal product of labor / Wage rate) = (Marginal product of capital / Rental rate)

Here,

Marginal product of labor / Wage rate = 500/100 = 5

Marginal product of capital / Rental rate = 1,000/200 = 5

Since (Marginal product of labor / Wage rate) = (Marginal product of capital / Rental rate), firm is employing cost-minimizing input combination.